# Rhapsody: A Dataset for Highlight Detection in Podcasts

### v1

@@ 5 Related Work
The podcast highlight detection task also falls under fine-grained information access tasks, as it focuses on identifying the most engaging segments within a podcast episode. Clifton et al. (2020) introduced 100,000 Podcasts, the largest podcast corpus to date, along with their case study on spoken passage retrieval. Using this corpus, Jones et al. (2020) and Karlgren et al. (2021) further ran a track on podcast segment retrieval and received submissions from the public, while providing tasks, human annotations, and evaluation measures. Recently Ghazimatin et al. (2024) proposed PODTILE, which automatically generates chapters for a podcast episode, given its metadata and transcript. However, all of these works relied on human annotation, thus limited to smaller scale study.
@@ 6 Conclusion
In this work, we introduce the task of podcast highlight detection and present Rhapsody, a dataset of 13K podcast episodes and binary highlight labels derived from YouTube’s replay graph feature. We benchmark the performance of different approaches with five automatic evaluation metrics. Our benchmark results demonstrate that even state-of-the-art LLMs, such as GPT-4o, struggle with this task, performing on par with simple heuristics. In contrast, our fine-tuned 1B Llama model significantly outperforms all the other approaches. However, the absolute performance of the best model remains low, underscoring the challenges of detecting meaningful highlights in podcast episodes. Future work can explore improved modeling approaches, such as incorporating different audio or contextual cues, to further enhance podcast highlight detection. Our system relies on separate components for transcription, summarization, audio feature extraction, and highlight prediction, which may introduce compounding errors. Future work can also explore an end-to-end approach that jointly optimizes these steps to enhance overall performance.
@@ A.1 Data Filtering
    "defaultLanguage" in channel_md["snippet"] and channel_md["snippet"]["defaultL%anguage"] != "en"
@@
    "podcastStatus" in playlist_md["status"] and playlist_md["status"]["podcastSta%tus"] == "enabled"
@@
    all([word not in playlist_md["snippet"]["title"] for word in ["music", "reacti%on", "review", "breakdown"]])
@@
    isEnglish(’ ’.join([playlist_md["snippet"]["title"], playlist_md["snippet"]["c%hannelTitle"], playlist_md["snippet"]["description"]]))
@@ B.1 Transcript Segment Summarization
[…]
3:for $k=1$ to $10$ do
@@ B.2 Audio Feature Extraction
We use an off-the-shelf emotion recognition model, audeering/wav2vec2-large-robust-12-ft-emotion-msp-dim, for audio feature extraction. The model produces a 1024-dimensional embedding from an input audio file. To obtain a [100, 1024] matrix per episode, we first load the raw audio at a 16, [^0] Hz sample rate, split it into 100 segments, and process each through the model. These embeddings can be further used to predict valence, dominance, and arousal scores with their pretrained regression head. As shown in Figure 8, there is a slight but obvious difference in all dimensions. For example, the average score is slightly higher for highlighted segments across all dimensions.
@@ Appendix D Model Implementation Details
The model is trained using a batch size of 1 and a gradient accumulation step of 8. QLoRA is applied with a dropout rate of 0.3, a rank of 8, and an alpha value of 16. The optimizer used is AdamW with a weight decay of 0.01 and a learning rate of 0.0003. We train the model with 4 NVIDIA RTX8000 GPUs. The model outputs the indices in the graph above a certain threshold as predictions. We use the prediction threshold of 0.3 and 0.425 for text-only and text + audio settings, respectively. These values are determined by the authors’ manual tuning on the validation set. We take checkpoints with the best average metrics on the validation set.

### v2

@@ 5 Related Work
The podcast highlight detection task also falls under fine-grained information access tasks, as it focuses on identifying the most engaging segments within a podcast episode. Clifton et al. (2020) introduced 100,000 Podcasts, comprising nearly 60,000 hours of speech, along with their case study on spoken passage retrieval. Using this corpus, Jones et al. (2020) and Karlgren et al. (2021) further ran a track on podcast segment retrieval and received submissions from the public, while providing tasks, human annotations, and evaluation measures. Recently Ghazimatin et al. (2024) proposed PODTILE, which automatically generates chapters for a podcast episode, given its metadata and transcript. However, all of these works relied on human annotation, thus limited to smaller scale study.
@@ 6 Conclusion
In this work, we introduce the task of podcast highlight detection and present Rhapsody, a dataset of 13K podcast episodes and binary highlight labels derived from YouTube’s replay graph feature. We benchmark the performance of different approaches with five automatic evaluation metrics. Our benchmark results demonstrate that even state-of-the-art LLMs, such as GPT-4o, struggle with this task, performing on par with simple heuristics. In contrast, our fine-tuned 1B Llama model significantly outperforms all the other approaches. However, the absolute performance of the best model remains low, underscoring the challenges of detecting meaningful highlights in podcast episodes.
Future work can explore improved modeling and inference methods, such as incorporating rich audio or contextual cues to further enhance podcast highlight detection. Our system relies on separate components for transcription, summarization, audio feature extraction, and highlight prediction, which may introduce compounding errors. Joint optimization in an end-to-end framework or improved input processing may mitigate this. Additionally, the structure of replay graphs itself and adaptive strategies for different podcast genres are another promising direction.
@@ A.1 Data Filtering
    "defaultLanguage" in channel_md["snippet"] and channel_md["snippet"]["defaultLanguage"] != "en"
@@
    "podcastStatus" in playlist_md["status"] and playlist_md["status"]["podcastStatus"] == "enabled"
@@
    all([word not in playlist_md["snippet"]["title"] for word in ["music", "reaction", "review", "breakdown"]])
@@
    isEnglish(’ ’.join([playlist_md["snippet"]["title"], playlist_md["snippet"]["channelTitle"], playlist_md["snippet"]["description"]]))
@@ B.1 Transcript Segment Summarization
2:Initialize: $S_{0}\leftarrow\text{``None.''}$, $\{\hat{t}_{i}\}_{i=1}^{100}\leftarrow$ uninitialized array of size 100
3:for $k=1$ to 1010 do
@@ B.2 Audio Feature Extraction
We use audeering/wav2vec2-large-robust-12-ft-emotion-msp-dim for DVA feature extraction, and facebook/hubert-base-ls960 for HuBERT feature extraction. The model produces a 1024-dimensional and 512-dimensional embedding from an input audio file, respectively. To obtain 100 embeddings per episode, we first load the raw audio at a 16, [^0] Hz sample rate, split it into 100 segments, and process each through the model.
Specifically, the DVA embeddings can be further used to predict valence, dominance, and arousal scores with their pretrained regression head. As shown in Figure 9, there is a slight but obvious difference in all dimensions. For example, the average score is slightly higher for highlighted segments across all dimensions.
@@ Appendix D Model Implementation Details
The model is trained using a batch size of 1 and a gradient accumulation step of 8. QLoRA is applied with a dropout rate of 0.3, a rank of 8, and an alpha value of 16. The optimizer used is AdamW with a weight decay of 0.01 and a learning rate of 0.0003. We train the model with 4 NVIDIA RTX8000 GPUs. The model outputs the indices in the graph above a certain threshold as predictions. We use the prediction threshold of 0.3, 0.425, and 0.375 for text-only, text + DVA, and text + HuBERT settings, respectively. These values are determined by the authors’ manual tuning on the validation set. We take checkpoints with the best average metrics on the validation set.